.     Let’s consider the old saying, “If we eat an apple a day it keeps the doctor away!”, with respect to risk assessment. We assume person would consume approximately 7300 kg of apples over a lifetime (80 years). If we assume in any given week, the maximum quantity ingested is 1.75 kg of apples. The typical concentration of Alar (a pesticide used on the apples) residue on the apples is 1.8 ppm (1.8 mg/kg) for all the apples ingested over a life time.

a.   A. What is the chronic daily intake (CDI) of Alar over a lifetime? Assume a bioavailability of 100 percent.
Answer #1

Total apple ingested in a week = 1.75 kg

Total pesticide ingested= 1.75*1.8= 3.15 mg per week

Pesticide ingested per day= 3.15/7= 0.45 mg per day

Total apple ingested in life time = 7300 kg

Total pesticide ingested in life time = 7300*1.8= 13140 mg = 13.14 grams